OpenStudy (anonymous):

From the following heats of combustion, CH3OH(l) + 3/2O2(g) → CO2(g) + 2H2O(l) ΔHorxn = –726.4 kJ/mol C(graphite) + O2(g) → CO2(g) ΔHorxn = –393.5 kJ/mol H2(g) + ½O2(g) → H2O(l) ΔHorxn = –285.8 kJ/mol Calculate the enthalpy of formation of methanol (CH3OH) from its elements. C(graphite) + 2H2(g) + ½O2(g) → CH3OH(l)

OpenStudy (anonymous):

I got this from a previous session but they didn't explain where the Carbon wen to in the one section. Number equations 1,2,3 Now, Invert EQN1 CO2(g) + 2H2O(l) → CH3OH(l) + 3/2O2(g) ΔHorxn = +726.4 kJ/mol Multiply EQN3 by 2 2H2(g) + O2(g) → 2H2O(l) ΔHorxn = 2*(–285.8 kJ/mol) = -571.+6 kJ/mol ADd equations CO2(g) + 2H2O(l) → CH3OH(l) + 3/2O2(g) ΔHorxn = +726.4 kJ/mol 2H2(g) + O2(g) → 2H2O(l) ΔHorxn = 2*(–285.8 kJ/mol) = -571.+6 kJ/mol CO2(g) + 2H2O(l) + 2H2(g) + O2(g) --> CH3OH(l) + 3/2O2(g) + 2H2O(l) ΔHorxn = +154.8 kJ/mol Cancel common terms CO2(g) + + 2H2(g) --> CH3OH(l) + 1/2O2(g) ΔHorxn = +154.8 kJ/mol Now, add EQN3 C(graphite) + O2(g) → CO2(g) ΔHorxn = –393.5 kJ/mol add both equations CO2(g) + + 2H2(g) --> CH3OH(l) + 1/2O2(g) ΔHorxn = +154.8 kJ/mol C(graphite) + O2(g) → CO2(g) ΔHorxn = –393.5 kJ/mol CO2(g) + + 2H2(g) + C(graphite) + O2(g) --> CH3OH(l) + 1/2O2(g) +CO2(g) ΔHorxn = -238.7 kJ/mol cancel terms 2H2(g) + C(graphite) + 1/2O2(g) --> CH3OH(l) ΔHorxn = -238.7 kJ/mol NOTE that Hf of C = 0 so... 2H2(g) + 1/2O2(g) --> CH3OH(l) ΔHorxn = -238.7 kJ/mol Now here is the question I have what happened to the graphite in multiplying equations 3 and 2.